### v1

# Theoretical Modeling of LLM Self-Improvement Training Dynamics Through Solver-Verifier Gap

Abstract
Self-improvement is among the most prominent techniques within the realm of large language models (LLM), aiming to enhance the LLM performance without relying on external data. Despite its significance, generally how LLM performances evolve during the self-improvement process remains underexplored. In this paper, we theoretically model the training dynamics of self-improvement via the concept of solver-verifier gap. This is inspired by the conjecture that the performance enhancement of self-improvement stems from the gap between LLM’s solver capability and verifier capability. Based on the theoretical framework, we further introduce how to predict the ultimate power of self-improvement using only information from the first few training epochs. We empirically validate the effectiveness of the theoretical model on various LLMs and datasets. Beyond self-improvement, we extend our analysis to investigate how external data influences these dynamics within the framework. Notably, we find that under limited external data regimes, such external data can be utilized at any stage without significantly affecting final performances, which accords with the empirical observations.
{NoHyper}**footnotetext: Equal contributions.
1 Introduction
Large language models (LLMs) have emerged as one of the most pivotal frontiers in artificial intelligence, propelling the development of diverse applications such as chatbots (Brown et al., 2020), mathematical reasoning (Wei et al., 2022; Shao et al., 2024), and robotics (Wu et al., 2023). Despite their remarkable success, the training of LLMs typically necessitates massive data. In practice, data collection often confronts significant challenges, and there are even concerns that available data sources could be depleted in the foreseeable future (Villalobos et al., 2024; Shen et al., 2025; Muennighoff et al., 2023; Wang et al., 2024a). This data bottleneck motivates researchers to explore alternative training strategies (Gao et al., 2020; Dong et al., 2024).
Among these strategies, a growing body of work focuses on training or fine-tuning LLMs using the data they generate, a process known as self-improvement (Bai et al., 2022; Huang et al., 2023; Wang et al., 2023; Pang et al., 2024). Self-improvement methodologies initiate with a pre-trained LLM, utilize the model to generate new data, and then fine-tune the model with the generated data. Empirical studies have shown that this approach yields promising results across various domains (Zelikman et al., 2022; Wang et al., 2023; Tian et al., 2024). However, the theoretical underpinnings of self-improvement remain under-explored (Song et al., 2025; Huang et al., 2025). Specifically, there is a lack of comprehensive theoretical models to explain its mechanisms, and insufficient evidence exists to fully understand the training dynamics involved in this process.
Figure 1: Overview of the theoretical framework on self-improvement and cross-improvement.
In this paper, we theoretically model the training dynamics of LLM self-improvement, inspired by the conjecture that self-improvement capability arises from the gap between an LLM’s solver capability and verifier capability (Song et al., 2025). Specifically, we define these two capabilities as follows:
  • Solver capability $U_{s}(t)$: The quality of responses directed generated from LLM;
  • Verifier capability $U_{v}(t)$: The quality of responses generated from LLM and evaluated by itself;
In practice, different metrics can be used to quantify these capabilities. For tasks with ground-truth labels, the $0-1$ training loss can serve as a direct measure. For tasks without ground truth, uncertainty quantification can be adopted, as it correlates strongly with model capability (Huang et al., 2025). Based on the above discussions, we model the training dynamics of LLM self-improvement using the following coupled differential equations, inspired by the term potential energy in physics (Rankine, 1853):
$\begin{split}&\frac{dU_{s}(t)}{dt}=-\alpha E(t),\quad\frac{dU_{v}(t)}{dt}=-\beta E(t),\end{split}$ (1)
where $\alpha,\beta$ denote the coefficients, $t$ denotes the epoch, $U_{s}(t)$ denotes the solver capability, $U_{v}(t)$ denotes the verifier capability, and $E(t)$ denotes the capability gap related to $U_{s}(t)-U_{v}(t)$. We omit the initial conditions for simplicity. Under such a framework, the resulting dynamics would be
$\begin{split}&U_{s}(t)=\alpha^{\prime}e^{-k(\alpha-\beta)t}+U_{s,\infty},\quad U_{v}(t)=\beta^{\prime}e^{-k(\alpha-\beta)t}+U_{v,\infty},\end{split}$ (2)
where $\alpha^{\prime},\beta^{\prime},k$ denotes the constants with detailed formulations in Section 3.2, and $U_{s,\infty},U_{v,\infty}$ represent the solver capability and the verifier capability at convergence, respectively. Along the trajectory, both capabilities follow exponential convergence laws according to the framework. Besides, we specifically focus on the solver’s ultimate capability $U_{s,\infty}=\frac{1}{\alpha-\beta}(\alpha U_{v,0}-\beta U_{s,0}+\alpha\frac{b}{k})$. This result reveals that the solver’s ultimate capability relates to both the initial capability $U_{s,0},U_{v,0}$ and the coefficient $\alpha,\beta$. Therefore, the ultimate capability might be improved given a larger verifier-solver gap at initialization, in accordance with our insights. Detailed derivations are provided in Section 3.2.
From the experimental perspective, we find in Section 4 that such theoretical modeling demonstrates strong practical efficacy. Specifically, experimental results across various models and datasets in Figure 2 reveal that the capability dynamics indeed follow an exponential law, as implied by the theoretical framework in Section 3. Besides, we observe in Figure 2 that the verifier capability consistently outperforms the solver throughout the self-improvement process, which might be empirical evidence of the solver-verifier gap’s crucial role in driving capability improvement. To further investigate the role of the solver-verifier gap, we conduct experiments in Section 4.3 on multiple LLMs demonstrating that the gap generally happens in practice. These findings hold across varying sample sizes (Figure 4) and even under cross-evaluation scenarios (Figure 5).
We further analyze in Section 5 the application of this framework in cross-improvement, a potential approach to enhance the ultimate capability of self-improvement. Cross-improvement in this paper refers to the utilization of external data in the verification step (details in Figure 7). Within the theoretical framework, we contend that cross-improvement outperforms self-improvement due to the enhanced verification capabilities. We then derive the training dynamics under the cross-improvement regimes, with enhanced verification capabilities compared to self-improvement. This analysis can further assist in answering the question: given the limited amount of external data, how should one allocate these external data during the cross-improvement training process? Our theoretical analyses demonstrate that the timing of using external data is not crucial; rather, the utilization of such data genuinely enhances the training process. Therefore, external data can be incorporated at any stage as desired. Experiments in Section 5.2 on the cross-improvement validate the theoretical findings. Our contributions can be summarized as follows:
  • Theoretical framework: We detail our theoretical framework on the dynamics of self-improvement in Section 3, inspired by the potential energy framework based on the solver-verifier capability gap in Equation 8. Theoretical derivations in Equation 13 imply an exponential law for the model capability, and we further establish the extreme capability based on the framework.
  • Experiments: We conduct experiments on self-improvement to verify the theoretical framework in Section 4 across multiple models, datasets, and verification methods. Our empirical observations indicate that (1) the uncertainty/accuracy during the self-improvement process indeed follows an exponential law with an extreme capability (Figure 2) as implied by the theoretical framework; (2) the solver-verifier gap generally happens in practice across different regimes (Figure 4 and Figure 5), validating the utility of the gap as the potential energy;
  • Cross-improvement: We further investigate in Section 5 the cross-improvement under the above framework, where limited external data is provided during the training process. We contend in Section 5.1 that cross-improvement might improve the verifier capability, thus improving the extreme capability of self-improvement. Empirically, we observe in Section 5.2 that if the verifier capability is improved by the external data, the performances of LLM are enhanced, and vice versa (Table 1). This accords with the theoretical findings.
2 Related Work
Self-Improvement. Self-improvement aims to enhance model performance without relying on external information (Huang et al., 2023; Wang et al., 2023; Bai et al., 2022; Pang et al., 2024). Self-improvement is of significant importance as it enables models to adapt and evolve autonomously, thereby facilitating their effectiveness across a wide range of real-world scenarios such as reasoning (Zelikman et al., 2022; Peng et al., 2024; Huang et al., 2023), alignment (Wang et al., 2023, 2024b; Ding et al., 2024), and planning (Tian et al., 2024). In this paper, we consider a branch of self-improvement that fine-tunes with the output of LLM (Amini et al., ; Sessa et al., 2024; Gui et al., 2024; Pace et al., ; Ouyang et al., 2022; Rafailov et al., 2023). Alternative approaches to self-improvement include self-distillation (Buciluǎ et al., 2006; Hinton et al., 2015; Zhang et al., 2019) which involves transferring knowledge from a larger, more complex model to a smaller one, and self-correction (Kumar et al., 2024; Liu et al., 2024) where the model identifies and rectifies its own errors, etc. A body of research has also explored the potential negative consequences of self-improvement, including degradation issues (Bertrand et al., 2024; Gerstgrasser et al., 2024) and failures on out-of-domain reasoning tasks (Yuan et al., 2025).
Theoretical Understandings on Self-Improvement. Theoretical insights into self-improvement could potentially enhance comprehension, thereby making self-improvement more reliable (Yampolskiy, 2015). Previous works have theoretically studied self-improvement via self-distillation techniques, providing convergence rates for linear models (Mobahi et al., 2020; Frei et al., 2022; Das and Sanghavi, 2023; Pareek et al., 2024), neural networks (Allen-Zhu and Li, 2023), and general models (Boix-Adsera, 2024). In the realm of LLM, several works have theoretically explored self-improvement with in-context alignment (Wang et al., 2024c), reinforcement learning (Talvitie, 2017; Choi et al., 2024; Gandhi et al., 2025), meta learning (Kirsch and Schmidhuber, 2022), and diffusion models (Fu et al., 2024). Nevertheless, the theoretical understanding of self-improvement in the context of LLM training dynamics still remains underexplored.
Most relevant to our work is Song et al. (2025) and Huang et al. (2025). Song et al. (2025) posits that the key to self-improvement lies in the generation-verification gap and further examines the relationship between this gap and pre-training flops. Huang et al. (2025) further posits that the improvement stems from a sharpening mechanism, in which the verification step sharpens the model performance on the high-quality sequences. Our paper draws inspiration from Song et al. (2025); Huang et al. (2025), as we employ the concept of a solver-verification sharpening gap in the theoretical analysis, and the training policy in this paper follows Huang et al. (2025). However, different from Song et al. (2025); Huang et al. (2025), our research primarily centers on developing the self-improvement dynamics based on the solver-verification sharpening gap. Additionally, we delve deeper into understanding how cross-improvement works within this theoretical framework.
Cross-Improvement. Besides self-improvement approaches, a branch of papers focuses on enhancing the capabilities of LLM through external data, namely, cross-improvement. One of the most frequently utilized sources of external data is human-annotated data (Ouyang et al., 2022; Lightman et al., 2024; Borchers et al., 2025). Despite its utility, the collection of such data is extremely resource-intensive. Moreover, relying solely on human-annotated data restricts the potential for LLM to surpass human performance. Another potential source of external data stems from stronger models (Ho et al., 2023; Chang et al., ; Lee et al., 2024), while access to these stronger models often presents significant challenges. Our paper also considers the scenario of cross-improvement that leverages a limited number of tokens from stronger models.
3 Theoretical Modeling of Self-Improvement Training Dynamics
This section presents a theoretical framework for sketching training dynamics in LLM self-improvement. We start by introducing necessary notations of self-improvement in Section 3.1. We then theoretically model the self-improvement dynamics in Section 3.2.
3.1 Preliminaries
This section introduces the basic notations and definitions, as well as the definitions of solver capability and verifier capability. We start from the basic notations on data and models.
Notations. Let $(x,y)$ denote a prompt-response pair, with $y^{[k]}$ denoting the $k$-th token and $y^{[1:k]}$ denoting the first $k$ tokens. We use $L(y)$ to represent the length of $y$. Let $\pi_{f}(y|x)$ denote the probability that a model $f$ generates a response $y$ given a prompt $x$, where $\pi_{f}(y|x)$ can be split with the auto-regressive structure of the response $\pi_{f}(y|x)=\prod_{k=1}^{L(y)}\pi_{f}(y^{[k]}|y^{[1:k-1]},x)$. We denote the best response as $y^{*}$, that is, the ground truth response. Ideally, a model’s performance could be measured by its loss relative to this ground truth, for instance, $L_{f}(\hat{y})=\|y^{*}-\hat{y}\|$. However, as not all tasks have an accessible ground truth, a different metric is required. Therefore, we also use the uncertainty metric following Huang et al. (2025) in our framework. We define the uncertainty for a response $\hat{y}$ given its prompt $x$ and a model $f$ as its negative log-likelihood:
$U_{f}(\hat{y})=-\log\pi_{f}(\hat{y}|x).$ (3)
This uncertainty, $U_{f}(\hat{y})$, will be our primary measure throughout the paper. The model’s capability is inversely related to the uncertainty: a lower uncertainty signifies a higher capability.
Solver. The solver is regarded as the model capability to return responses with low uncertainty. Therefore, for each prompt $x_{i}$, we sample one response $\hat{y}_{i}(t)$ to represent the solver solution. That is,
$\hat{y}_{i}\sim\pi_{f}(\cdot|x).$ (4)
Note that we only draw one response for each prompt due to calculation efficiency. An alternative solution is to generate multiple responses and use the whole distribution, but the two policies are similar since we already take average operations over (i.i.d.) prompts.
Verifier. We use LLM itself as the verifier. For each prompt $x_{i}$, we first sample $N$ responses based on the LLM output $\hat{y}_{i,1},\cdots,\hat{y}_{i,N}\sim\pi_{f}(\cdot|x_{i})$. We then ask the LLM to evaluate these responses with a score $s(\hat{y}_{i,j})\in[0,1]$. The Best-of-N (BoN) response is then defined as
$\hat{y}_{i}^{\text{BoN}}=\mathop{\mathrm{argmin}}_{\{\hat{y}_{i,j}:s(\hat{y}_{i,j})\geq\sigma\}}\frac{1}{L(\hat{y}_{i,j})}U_{f}(\hat{y}_{i,j}|x_{i}),$ (5)
where $\sigma$ denotes the threshold parameter, and we use $1/L(\hat{y}_{i,j})$ as a regularizer to discourage those short responses. The BoN solution first eliminates those solutions with small scores $s(\hat{y}_{i,j})$, and then finds the solution with the best capability. We deploy such mixed strategy to enhance the computational stability; as a comparison, strategy with only score $s(\hat{y}_{i,j})$ might make $\hat{y}_{i}^{\text{BoN}}$ has large variance. Obviously, the BoN solution merges the LLM verifier capability (via the score $s(\cdot)$) and the LLM output (via the uncertainty $U_{f}(\cdot)$). Therefore, the verifier capability can be calculated based on the uncertainty of the BoN solution. We finally remark that our paper uses a slightly different BoN policy compared to Huang et al. (2025) where they do not eliminate those responses with low scores, since we want to include more verification capability in the BoN response.
Solver and Verifier Uncertainty Metrics. Based on the above discussion, we use the average uncertainty of LLM response $\hat{y}$ to represent the solver capacity and use the average uncertainty of BoN response $\hat{y}^{BoN}$ to represent the verification capability. Therefore, we define the solver uncertainty $U_{s}(t)$ and verifier uncertainty $U_{v}(t)$ as
$\begin{split}&U_{s}(t)\triangleq\sum_{i=1}^{n}U_{f}(\hat{y}_{i}(t))=-\frac{1}{n}\sum_{i=1}^{n}\log\ \pi_{f}(\hat{y}_{i}(t)|x_{i}),\\ &U_{v}(t)\triangleq\sum_{i=1}^{n}U_{f}(\hat{y}^{\text{BoN}}_{i}(t))=-\frac{1}{n}\sum_{i=1}^{n}\log\ \pi_{f}(\hat{y}_{i}^{BoN}(t)|x_{i}),\end{split}$ (6)
where $\hat{y}_{i}(t)$ denotes the LLM output. In our framework, uncertainty metrics serve as inverse metrics of capability: a lower uncertainty value ($U_{s}$ or $U_{v}$) indicates a higher corresponding capability. Note that both $U_{s}(t)$ and $U_{v}(t)$ contains randomness, since $\hat{y}_{i}(t)$ is randomly generated by LLM, and the score $s(\cdot)$ in $\hat{y}^{\text{BoN}}_{i}(t)$ is also randomly generated by LLM. However, since the prompts $x_{i}$ are independent, the randomness could be controlled when the number of prompts $n$ is large.
Self-Improvement. We deploy self-improvement based on the above solver-verifier framework, similar to Huang et al. (2025). Notably, the verifier is slightly different, since we want to include more verification information. Overall, we first generate responses from the LLM. We then choose BoN response based on Equation 5. The optimization objective is to minimize the average uncertainty of BoN responses. The loss function $L_{t}(f)$ for a training step $t$ with function $f$ is defined as the verifier uncertainty $U_{v}(t)$:
$L_{t}(f)\triangleq U_{v}(t)=-\frac{1}{n}\sum_{i=1}^{n}\log\ \pi_{f}(\hat{y}_{i}^{\text{BoN}}(t)|x_{i}).$ (7)
By minimizing Equation 7, we steer the model to increase the likelihood of generating high-quality responses, effectively improving its solver capability. We summarize the one-step self-improvement algorithm in Algorithm 1.
Algorithm 1 Self Improvement (One Step)
0:  Prompts set $\mathcal{X}$:$\{x_{1},x_{2},\cdots,x_{n}\}$, Current model $f$, Sample Size $N$, Threshold $\sigma$
1:  $\mathcal{Y}^{\text{BoN}}\leftarrow\emptyset$, $\mathcal{U}^{\text{BoN}}\leftarrow\emptyset$;
2:  for each prompt $x_{i}\in\mathcal{X}$ do
3:     $C_{x_{i}}\leftarrow\emptyset$;
4:     for $j\leftarrow 1$ to $N$ do
5:        Generate response $\hat{y}_{i,j}\sim\pi_{f}(\cdot|x_{i})$;
6:        Ask model to evaluate $\hat{y}_{i,j}$ and return a score $s(\hat{y}_{i,j})$;
7:        if $s(\hat{y}_{i,j})\geq\sigma$ then
8:           Append $\hat{y}_{i,j}$ to $C_{x_{i}}$;
9:        end if
10:     end for
11:     $\hat{y}_{i}^{\text{BoN}}\leftarrow\mathop{\mathrm{argmin}}_{C_{x_{i}}}\frac{1}{L(\hat{y}_{i,j})}U_{f}(\hat{y}_{i,j}|x_{i})$;
12:     Append $\hat{y}_{i}^{\text{BoN}}$ to $\mathcal{Y}^{\text{BoN}}$;
13:     Append $U_{f}(\hat{y}_{i}^{\text{BoN}})$ to $\mathcal{U}^{\text{BoN}}$;
14:  end for
15:  $\text{Uncertainty}\leftarrow\frac{1}{|\mathcal{U}^{\text{BoN}}|}\sum_{u\in\mathcal{U}^{\text{BoN}}}u$;
16:  $\hat{f}\leftarrow\operatorname{AdamW}(f,\text{Uncertainty})$;
16:  Self-improved model $\hat{f}$
3.2 Self-Improvement Dynamics
In this section, we aim to analyze the dynamics of self-improvement. Our techniques are inspired by the concept of potential energy, a widely used concept in physics (Rankine, 1853). Following Huang et al. (2025) and Song et al. (2025), we argue that the self-improvement comes from the Capability Gap $G(\cdot)$, defined as the gap between $U_{s}(t)$ and $U_{v}(t)$, namely,
$G(t)\triangleq U_{s}(t)-U_{v}(t)=-\frac{1}{n}\sum_{i=1}^{n}\log\ \frac{\pi_{f}(\hat{y}_{i}(t)|x_{i})}{\pi_{f}(\hat{y}_{i}^{BoN}(t)|x_{i})}.$ (8)
We assume that the change of the solver and verifier capability (Equation 6) is driven by a gap potential energy $E(t)$, assumed as a linear function of the Capability Gap $G(t)$ (Equation 8), namely
$E(t)=kG(t)-b,$ (9)
where $k,b$ denote the linear coefficient and bias. Experiments in Section 4 show that assuming potential energy in a linear form (as in Equation 9) is enough to explain empirical phenomena in self-improvement training. In this paper, we simply use uncertainty to evaluate the response quality, leading the capability gap defined with the uncertainty, as discussed in Section 3.1. We assume that the solver capability and the verifier capability both increase during the process of self-improvement, which is widely observed in the related works (Song et al., 2025). With the analysis above, following the concept of potential energy, the theoretical framework starts with the following assumptions:
$U_{s}(t)|_{t=0}=U_{s,0},$ $\quad U_{v}(t)|_{t=0}=U_{v,0},$ (10)
$\frac{dU_{s}(t)}{dt}=-\alpha E(t),$ $\quad\frac{dU_{v}(t)}{dt}=-\beta E(t),$ (11)
where $\alpha,\beta\geq 0$ are coefficients related to the decreasing rate of $U_{s}(t),U_{v}(t)$. In this framework, Equation 10 represents the initial conditions. Since for LLM, the verification capability usually outperforms the solver capability in real-world applications, we assume that $U_{s,0}>U_{v,0}$quation 11 represents the iterative conditions, where we assume that $\alpha>\beta$, indicating that the solver capability increases faster than the verifier. Based on the above assumptions, we derive the following dynamics:
$E(t)=k\gamma e^{-k(\alpha-\beta)t},$ $\quad G(t)=\gamma e^{-k(\alpha-\beta)t}+G_{\infty},$ (12)
$U_{s}(t)=\alpha^{\prime}e^{-k(\alpha-\beta)t}+U_{s,\infty},$ $\quad U_{v}(t)=\beta^{\prime}e^{-k(\alpha-\beta)t}+U_{v,\infty},$ (13)
where $\gamma=U_{s,0}-U_{v,0}-\frac{b}{k}$, $\alpha^{\prime}=\frac{\alpha}{\alpha-\beta}(U_{s,0}-U_{v,0}-\frac{b}{k})$, $\beta^{\prime}=\frac{\beta}{\alpha-\beta}(U_{s,0}-U_{v,0}-\frac{b}{k})$ represent coefficients, and $G_{\infty}=\frac{b}{k}$, $U_{s,\infty}=U_{s,0}-\alpha^{\prime}$, $U_{v,\infty}=U_{v,0}-\beta^{\prime}$ represent the capability at convergence. We demonstrate from Equation 13 that $U_{s}$ and $U_{v}$ both decrease to an inherent uncertainty level obeying exponential law during self-improvement with $U_{s}$ decreasing much faster. The dynamics also indicate that (i) the gap potential energy $E(t)$ drives the solver capability stronger; (ii) the change of solver capability and verifier capability slows down as the gap decreases; and (iii) the capability gap might not necessarily converge to zero during the training process. We empirically validate our theoretical framework in Section 4.
4 Experiment
In this section, we conduct experiments on self-improvement to verify the theoretical framework. We first introduce experimental setups in Section 4.1. We then present the main experimental results in Section 4.2, sketching the dynamics of the self-improvement process. We finally explore the differences between the solver capability and the verifier capability, showing that the solver-verifier gap indeed generally happens in practice.
Figure 2: Accuracy and uncertainty during the self-improvement of the Phi4-mini model on the Math and GSM8k datasets using the QE method. The experimental results show that the accuracy increases during self-improvement process while the uncertainty decreases.
4.1 Setup
This section details the methods, models, datasets, and key parameters employed in our experiments. We consider the following verification methods: TrueFalse (TF): The solver generates $N$ responses, denoted $\hat{y}_{i,1},\cdots,\hat{y}_{i,N}$, for a prompt $x_{i}$. The verifier is then tasked with answering whether each response $\hat{y}_{i,j}$ is correct. If the verifier deems a response $\hat{y}_{i,j}$ correct, its score $s(\hat{y}_{i,j})$ is set to $1$; otherwise, it is set to 00; Quality Evaluation (QE): The solver generates $N$ responses, $\hat{y}_{i,1},\cdots,\hat{y}_{i,N}$, for a prompt $x_{i}$. The verifier then assigns a continuous score $s(\hat{y}_{i,j})$ between 0 and 1 to each response based on its quality. A score of $s(\hat{y}_{i,j})=0$ indicates a completely incorrect answer, while $s(\hat{y}_{i,j})=1$ indicates a completely correct answer.
We utilize two model families in our study: (a) Phi models: From the Phi family (Abdin et al., 2024), we use Phi4-Mini, Phi3.5-Mini, and Phi3-mini; (b) Llama models: We use Llama3.2-3B (Grattafiori et al., 2024) and Llama3.1-8B. Our experiments focus on the models’ mathematical problem-solving capabilities. Accordingly, we employ two representative datasets: GSM8k (Cobbe et al., 2021) and Math (Hendrycks et al., 2021). To maintain response diversity, we set the temperature to $1$, maximum length of $512$ tokens, threshold $\sigma=0.5$, and sample size $N=16$. Zero-shot prompting is used for all models and datasets. If the verifier cannot provide a definitive evaluation for an answer, we treat it as a neutral response and assign a score of $0.5$. Experimental details are provided in Appendix A.
4.2 Dynamics of Self-Improvement
This section sketches the dynamics of self-improvement using AI-generated feedback for supervised fine-tuning (SFT), aiming to verify the theoretical framework proposed in Section 3.2.
Setups. We tested the Phi4-mini, Phi3.5-mini, Phi3-mini, and Llama3.2-3B models. We applied Low-Rank Adaptation (LoRA) (Hu et al., 2022), which significantly reduces the number of updatable parameters, thereby enhancing SFT efficiency. The chosen hyperparameters are detailed in Section A.1. To improve training efficiency, we employed mini-batch gradient descent with a batch size of $256$. For each model-task pair, we conducted training for $10$ epochs, saving a checkpoint after processing half of the training data. We test the solver accuracy and the verifier accuracy, defined as the accuracy of the response and the BoN response, respectively.
Results. We present the results of self-improvement on the Phi4-mini model using the QE method, as depicted in Figure 2. The empirical evidence indicates a consistent enhancement in the accuracy of both the solver and the verifier during the self-improvement process, coupled with a concurrent reduction in their respective uncertainties. Furthermore, a narrowing of the gap $G(t)$ between the solver and verifier is evident. Results for other models and methods are presented in Section A.2.
Validation. To validate our theoretical framework and enable the prediction of the model’s final performance, we fit an exponential model to the uncertainty from $10$ self-improvement epochs. Figure 3 presents the results for the Phi4-mini model in four different settings (Math / GSM8K datasets with QE / TF metrics). The plots illustrate the evolution of three key metrics: solver uncertainty, verifier uncertainty, and the uncertainty gap. In all subfigures, the exponential model demonstrates a strong fit to the empirical data, with the coefficients of determination ($R^{2}$) exceeding $0.9$. This empirical evidence validates the exponential law proposed in our theoretical work.
Figure 3: Exponential trends of model uncertainty during self-improvement. The results illustrate the uncertainty associated with the Phi4-mini model’s solver and verifier, as well as their gap. The scatter points represent the measured data, while the solid lines are the best-fit curves to an exponential model. $R^{2}>0.9$ indicates that the evolution of these uncertainties is well-described by an exponential function.
Figure 4: Accuracy and uncertainty of Phi4-mini on Math and GSM8k with different sample size using TF and QE respectively. The results illustrate that the verifier perform better than the solver.
4.3 Verifier Outperform Solver
In this section, we evaluate the solver and verifier performance of LLM, aiming to validate the utility of the solver-verifier gap used in Section 3. Specifically, the experimental results verify that verifier capability outperforms solver capability consistently. To evaluate the comprehensive capability of the model in the training set and test set, we randomly sample $2,048$ instances of the training set and $1,024$ instances from the test set for each data set. Our evaluation is divided into two settings: self-evaluation and cross-evaluation. Self-evaluation employs the same model for both the solver and verifier roles, whereas cross-evaluation utilizes different models for the solver and verifier.
Self Evaluation. In this part, we compare the accuracy of solver and verifier on different models and datasets. Given that the verifier selects one response from $N$ candidates generated by the solver, the verifier’s performance is expected to improve as $N$ increases. For each model, we evaluate the accuracy and uncertainty of both the solver and the verifier, varying $N$ across the values ${2,4,8,16,32,64}$. As illustrated for the Phi4-mini model in Figure 4, an increase in $N$ corresponds to improved verifier accuracy and reduced uncertainty. This figure also shows that the verifier consistently outperforms the solver for Phi4-mini, achieving higher accuracy and lower uncertainty. Similar trends were observed across other tested models, with the possible exception of Phi4. Detailed results are presented in Section A.1.
Cross Evaluation. In self-evaluation, the same model serves as both solver and verifier. To better understand the relationship between solver and verifier capabilities, we also perform cross-evaluation, where one model acts as the solver and a different model acts as the verifier. These cross-evaluations utilize models with the QE method on both the MATH and GSM8k datasets. Furthermore, we set $N=16$ because, as demonstrated in Figure 4, accuracy does not improve significantly for $N>16$. The results of these evaluations are presented in Figure 5. We observe that when a fixed model serves as the solver, the verifier’s performance generally surpasses that of the solver, even when a different model is employed as the verifier.
These experiments indicate that the verifier typically outperforms the solver, revealing a consistent positive performance gap between the verifier and the solver across model-task pairs. This performance gap is considered a key driver of self-improvement dynamics.
Figure 5: Cross-evaluation using QE method with sample size $N=16$ on Math. For each solver model, sampled responses are verified by different models. Figure on the left illustrates the difference of accuracy while figure on the right shows the $10$ logarithms of the uncertainty difference.
4.4 Pass@K
In this section, we investigate the underlying reason for the limit of self-improvement, focusing on how a decrease in the model’s response diversity leads to a diminishing potential energy, thereby causing the model’s capability to plateau. Although the self-improvement paradigm shows great improvement in model capability, it iteratively leads to a performance plateau. To investigate the underlying cause of this saturation, we conduct the Pass@K experiment on the initial model and the self-improved model. Pass@K is a metric that calculates the proportion of prompts where at least one of the $K$ responses is correct. When $K$ is large, Pass@K could be used to measure the diversity of the solver. We present the result of QE method in Figure 6. We observe that when $K$ is small, Pass@K increases with the number of epochs of self-improvement, validating the self-improvement process. However, when $K$ is large, we observe a slight decrease in Pass@K, indicating that the diversity of the solver is reduced through self-improvement. The degradation in diversity is caused by the convergence to a certain response, which is a potential reason for the limit of self-improvement.
Figure 6: Pass@K with QE method for different $K$ at $t=0$, $t=5$ and $t=10$. Pass@K evaluates the diversity of model generation. The efficacy of the self-improvement process is demonstrated by an increase in the Pass@K metric for small values of $K$. Conversely, for large values of $K$, a decrease in Pass@K is observed. This phenomenon suggests that the diversity of generations decreases during the self-improvement process.
5 Discussions on Cross-Improvement
In this section, we discuss how to model cross-improvement within the framework described above. The key insight is that external data affects the theoretical framework through the verification capability. We first present the theoretical framework in Section 5.1, under the framework with limited external data. We then conduct experiments in Section 5.2 to validate the theoretical findings.
5.1 Theoretical Framework of Cross-Improvement
In this section, we present the theoretical framework of training dynamics of cross-improvement. We follow the notations in Section 3. Besides, assume that we have limited external data with size $N^{\prime}$. For example, we may acquire $N^{\prime}$ external data in total from a better LLM using API queries.
We focus on the allocation of the external data. Specifically, for each epoch $t$, only $\eta_{t}N^{\prime}$ prompts could use API queries to get (one) external data, with $\sum_{t=1}^{T}\eta_{t}=1$ where $T$ denotes the total training epochs. For those chosen prompts, we choose the external data as the BoN response; for those non-chosen prompts, we still choose the BoN data from the $N$ internal responses. Notably, as long as one prompt is chosen to use external data, it will always use the external data.
Cross-Data Effects. In the cross-improvement framework, the use of external data will influence the verifier capability $U_{v}(t)$, since we use a different definition of BoN which directly relates to the verifier capability. To model the effects, we assume the verifier capability after cross-improvement as
$U^{c}_{v}(t)=(1+\gamma\eta_{t})^{-1}U_{v}(t-1).$ (14)
This assumption is a simplification of the effect of cross-improvement, and is demonstrated by experimental observations. The parameter $\gamma$ represents the general effect of cross-improvement, while $\eta_{t}$ represents the ratio of external data used in epoch $t$. We assume that the effect $\gamma$ is time-invariant, without which one cannot estimate the parameter. Overall, in each step, the cross-data first boosts the verification capability, then it evolves following the mathematical framework in Section 3.2. We illustrate this procedure in Figure 7.
Solutions. We model the above framework on cross-improvement as
$U_{s}(t)|_{t=0}=U_{s,0},$ $\quad U_{v}(t)|_{t=0}=U_{v,0},$ (15)
$U_{s}^{c}(t)=U_{s}(t-1),$ $\quad U_{v}^{c}(t)=(1+\gamma\eta_{t})^{-1}U_{v}(t-1),$ (16)
$G^{c}(t)=U_{s}^{c}(t)-U_{v}^{c}(t),$ $\quad E(t)=kG^{c}(t)-b,$ (17)
$U_{s}(t)-U_{s}^{c}(t)=-\alpha E(t),$ $\quad U_{v}(t)-U_{v}^{c}(t)=-\beta E(t).$ (18)
Equation 15 represents the initial conditions, which are the same as Equation 10Equation 16 represents the effects of cross-improvement, where the solver capability after cross-improvement $U_{s}^{c}(t)$ remains unchanged, while the verifier capability increases as discussed in Equation 14. The current capability gap $G^{c}(t)$ is then defined as the gap between the solver capability and the current verifier capability. Equation 18 represents the iterative conditions of cross-improvement. Note that Equation 18 differs slightly from the self-improvement dynamics in Equation 11 in the following ways: (i) we employ the external-data-affected state of solver capability $U_{s}^{c}(t)$ and verifier capability $U_{v}^{c}(t)$ during the process; (ii) we adopt discrete iteration instead of using a SDE. This is because formulating the SDE becomes challenging under the changes described in (i). Based on the above formulation, we derive an approximate solution of the ultimate uncertainty:
$\mathbf{U}(T)\approx e^{-\mathbf{\Delta^{\prime}}}\mathbf{U}(0),$ (19)
$\mathbf{\Delta^{\prime}}=\begin{pmatrix}T-(1+\beta k)(T-\gamma\sum_{t=1}^{T}\eta_{t})&T\beta k&-T\beta b\\ -\alpha k(T-\gamma\sum_{t=1}^{T}\eta_{t})&T\alpha k&-T\alpha b\\ 0&0&0\end{pmatrix}.$ (20)
where $\mathbf{U}(T)$ denotes $\begin{bmatrix}U_{v,T},&U_{s,T},&1\end{bmatrix}^{\top}$ and $\mathbf{U}(0)$ denotes $\begin{bmatrix}U_{v,0},&U_{s,0},&1\end{bmatrix}^{\top}$. $U_{s}(T)$ is related only to the summation $\sum_{t=1}^{T}\eta_{t}$(instead of each $\eta_{t}$ at epoch $t$). Under the approximate solution, we come to the following conclusions:
  • For cross-improvement with $\sum_{t=1}^{T}\eta_{t}=1$, the approximate solution is around the same;
  • The cross-improvement with $\sum_{t=1}^{T}\eta_{t}=1$ outperforms self-improvement with $\sum_{t=1}^{T}\eta_{t}=0$ in terms of solver capability, when $\gamma>0$.
Detailed derivations of the above part are presented in Appendix B.
5.2 Experiments
Table 1: Solver accuracy with QE method on train set: raw data and relative improvements of strategies average (%). We note that the improvement of late strategy on Math dataset is slight, which may caused by the insufficient training on the external data.
Strategy Phi-4-mini-instruct Llama-3.2-3B-Instruct
Math GSM8k Math GSM8k
Initial 30.31 ($\pm$ 0.24) 73.42 ($\pm$ 0.33) 36.02 ($\pm$ 0.25) 63.10 ($\pm$ 0.58)
Baseline 43.87 ($\pm$ 0.36) 87.71 ($\pm$ 0.11) 49.16 ($\pm$ 0.73) 87.00 ($\pm$ 0.21)
Early 44.59 ($\pm$ 0.86) 87.72 ($\pm$ 0.24) 52.52 ($\pm$ 0.47) 86.35 ($\pm$ 0.32)
Uniform 45.21 ($\pm$ 0.50) 87.82 ($\pm$ 0.21) 52.73 ($\pm$ 0.34) 85.99 ($\pm$ 0.65)
Late 43.73 ($\pm$ 0.13) 87.71 ($\pm$ 0.21) 48.98 ($\pm$ 0.17) 85.07 ($\pm$ 0.33)
Max-Min 1.48 0.11 3.75 1.28
Avg 44.51 87.75 51.41 85.80
Avg vs Initial +14.20 +14.33 +15.39 +22.70
Avg vs Baseline +0.64 +0.04 +2.25 -1.20
Early($t=0$) vs Initial (Verifier) +5.87 +0.96 +0.97 -1.62
In this section, we perform cross-improvement experiments using different allocation strategies.  Figure 7 illustrates the process of cross-improvement. For these experiments, external data are generated by DeepSeek-V3. Specifically, we utilize DeepSeek-V3 responses for fine-tuning instead of BoN responses. We perform $10$ epochs of cross-improvement on the Phi4-mini and Llama3.2-3B models using the QE method. To facilitate the model’s learning from external data, all such data are introduced within the first $8$ training epochs; data introduced in an epoch remain available throughout $10$-epoch period. The total number of DeepSeek-V3 responses used is $3000$, and we test three allocation strategies:
Figure 7: A diagram of Cross Improvement
  • Early: All external data are introduced in the first epoch.
  • Uniform: An equal amount of new external data is introduced in each of the first eight epochs.
  • Late: All external data are introduced in the eighth epoch.
We present the results on the training data in Table 1. In this table, Initial represents the performance of the original model, while Baseline represents the results from self-improvement. We observe that the average solver accuracy for all three strategies is higher than the baseline on the MATH dataset. However, Phi4-mini shows a slight performance improvement on GSM8k, while Llama3.2-3B even exhibits a performance drop on this dataset.
To better understand why cross-improvement is less effective on GSM8k, we calculate the difference in verifier accuracy at $t=0$ between the Early strategy and the Initial model, where $t=0$ signifies that external data are incorporated but no training is performed. This difference is listed in the last row of Table 1. We find that this difference is small for Phi4-mini and negative for Llama3.2-3B. This implies that, on GSM8k, the verifier’s performance after the addition of external data is not significantly better than its performance without such data, and may even decrease. This observation indicates that a key factor influencing the effectiveness of cross-improvement is indeed whether the external model’s verifier capability significantly surpasses that of the original model.
Additionally, we also observe that the late strategy has only a slight improvement on the Math dataset. To determine the cause of this phenomenon, we conduct a $12$ epochs cross-improvement experiment on the Phi4-mini and Llama3.2-3B model and present the result in Table 2. The solver accuracy of late strategy rises to $45.83\%$ and $51.31\%$, respectively, indicating that the model has not fully learned the external data with late strategy in the original experiment. Furthermore, we observe that the performance differences in the solver accuracy between the three strategies are marginal. This suggests that the timing of external data integration does not significantly influence the effectiveness of the cross-improvement method.
Table 2: Solver accuracy of three strategies with QE method on train set for $12$ epoch-training. The result demonstrates that late strategy has similar result with early and uniform strategy after $12$-epoch training.
Strategy Phi-4-mini-instruct Llama-3.2-3B-Instruct
Initial 30.31 ($\pm$ 0.24) 36.02 ($\pm$ 0.25)
Baseline 43.87 ($\pm$ 0.36) 49.16 ($\pm$ 0.73)
Early 46.33 ($\pm$ 0.13) 53.47 ($\pm$ 0.33)
Uniform 46.56 ($\pm$ 0.15) 52.73 ($\pm$ 0.66)
Late 45.83 ($\pm$ 0.39) 51.31 ($\pm$ 0.62)
6 Conclusion
In this paper, we propose a theoretical framework to analyze the training dynamics of self-improvement via the solver-verifier gap. Experimental results on various datasets and models accord well with the theoretical findings. Besides, one may derive the theoretical limits based on the framework, which is closely related to the model’s verification capability. To break the limit, one may apply cross-improvement to enhance the model’s verification capability. Therefore, we further introduce the corresponding theoretical framework on the dynamics of cross-improvement under limited external data regimes, and find that the allocation of external data might have less influence on the final results. Experimental results verify the theoretical findings. In the end, we provide several potential future directions:
  • Although the experiments accord well with the theory, the mechanism behind the theoretical framework is still under-explored;
  • We assume a time-invariant property when analyzing the cross-improvement, which might be relaxed in future work.
  • Developing self-improvement algorithms based on our framework is a potential direction in the future.
  • External data can be used to fine-tune a well-self-improved model to further improve the performance of the model.
Appendix
This section provides the supplementary material. Appendix A provides experimental details and omitted experimental results. Appendix B provides omitted derivations of the theoretical models. Appendix C discuss the difference between our work and prior works, as well as the challenges and future direction for self-improvement.
Appendix A Experimental Details
In this section, we detail our experiment hyperparameters and present all results of our experiment. In Section A.1, we detail the hyperparameters chosen in our experiments. We provide all All of our experiments are run on 80G NVIDIA A800 GPUs.
A.1 Hyperparameters
In this section, we detail the hyperparameters in Table 3.
Table 3: Hyperparameters for SFT
Learning Rate Weight Decay LoRA Rank LoRA dropout Solver Temperature Verifier Temperature
1e-5 0.01 16 0.5 1 0.1
A.2 Omitted figures
In this section, we will provide figures omitted in Section 4.
Self-improvement In this part, we display figures omitted in Section 4Figure 8 illustrates the results of self-improvement on Phi3.5-mini, Phi3-mini and Llama3.2-3B model with QE method. We observe that most of model-task pairs have similar results with the Phi4-mini model except for several pairs. We also present the results with TF method in Figure 9. When self-improving the Phi3.5-mini and Phi3-mini models on Math data set, we observe that accuracy and uncertainty decrease simultaneously after several training epochs. The reason for this phenomenon might be LLM misleading by incorrect responses, as the BoN response may correspond to a incorrect answer with low uncertainty. The solution of this problem could be a future direction. Additionally, we note that the verifier accuracy of Llama3.2-3B on GSM8k decreases as training progresses and is lower than the solver accuracy after 8 epochs of training. One possible reason for this phenomenon is that we use length-regularized log-likelihood to obtain the BoN responses, so the BoN responses tend to be longer responses. Long responses are more likely to contain repeated content, which may reduce accuracy.
Self-evaluation This part we provide the results of self-evaluation for different sample size $N$Figure 10 illustrates the accuracy and uncertainty of Phi3.5-mini, Phi3-mini, Llama3.2-3B and Llama3.1-8B on Math and GSM8k with different sample size using TF and QE respectively, which shows that the verifier outperforms solver in all model-task pairs.
Cross-evaluation We present cross-evaluation results in Figure 11. It shows that the verifier always performs better than solver, although they use different LLMs.
A.3 Curve fitting
In this section, we present the fitted curves of Phi3.5-mini, Phi3-mini and Llama3.2-3B in Figure 12 and observe that all curves fitted well.
A.4 Pass@K
In this section, we present the Pass@K with TF method at different training epochs in Figure 13. We observe that Pass@K increases when K is small and decreases when K is large, which is similar to the results with the QE method.
(a) Phi3.5-mini
(b) Phi3-mini
(c) Llama3.2-3B
Figure 8: Accuracy and uncertainty during the self-improvement of the Phi3.5-mini, Phi3-mini and Llama3.2-3B on the Math and GSM8k datasets using the QE method.
(a) Phi4-mini
(b) Phi3.5-mini
(c) Phi3-mini
(d) Llama3.2-3B
Figure 9: Accuracy and uncertainty during the self-improvement of the Phi4-mini, Phi3.5-mini, Phi3-mini and Llama3.2-3B on the Math and GSM8k datasets using the TF method.
(a) Phi3.5-mini
(b) Phi3-mini
(c) Llama3.2-3B
(d) Llama3.1-8B
Figure 10: Accuracy and uncertainty of Phi3.5-mini, Phi3-mini, Llama3.2-3B and Llama3.1-8B on Math and GSM8k with different sample size using TF and QE respectively.
Figure 11: (a) Cross-evaluation using TF method with sample size N=16 on Math (b) Cross-evaluation using QE method with sample size N=16 on GSM8k (c) Cross-evaluation using TF method with sample size N=16 on GSM8k. For each solver model, sampled responses are verified by different models. Figures on the left illustrate the difference of accuracy while figures on the right show the 10 logarithms of the uncertainty difference.
(a) Phi3.5-mini
(b) Phi3-mini
(c) Llama3.2-3B
Figure 12: Uncertainty curve versus epochs of Phi4-mini, Phi3.5-mini, Phi3-mini and Llama3.2-3B models using QE and TF methods on Math and GSM8k datasets.
Figure 13: Pass@K with TF method at $t=0$, $t=5$ and $t=10$.
Appendix B Theoretical Derivation
In this section, we present the theoretical derivations of the conclusion presented in Section 5.1, from the dynamics conditions. The discussion and conclusion presented at the end of Section 5.1 stay the same.
According to Section 5.1, the dynamics conditions are formulated as:
$U_{s}(t)|_{t=0}=U_{s,0},$ $\quad U_{v}(t)|_{t=0}=U_{v,0},$ (21)
$U_{s}^{c}(t)=U_{s}(t-1),$ $\quad U_{v}^{c}(t)=(1+\gamma\eta_{t})^{-1}U_{v}(t-1),$ (22)
$G^{c}(t)=U_{s}^{c}(t)-U_{v}^{c}(t),$ $\quad E(t)=kG^{c}(t)-b,$ (23)
$U_{s}(t)-U_{s}^{c}(t)=-\alpha E(t),$ $\quad U_{v}(t)-U_{v}^{c}(t)=-\beta E(t).$ (24)
Denote $\mathbf{U}(t)$ as vector $\begin{bmatrix}U_{v}(t),&U_{s}(t),&1\end{bmatrix}^{\top}$, $\mathbf{U^{c}}(t)$ as vector $\begin{bmatrix}U_{v}^{c}(t),&U_{s}^{c}(t),&1\end{bmatrix}^{\top}$, with the following relationship holding true:
$\mathbf{U^{c}}(t)=\begin{pmatrix}\frac{1}{1+\gamma\eta_{t}}&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}\mathbf{U}(t-1).$ (25)
With these notations, we can derive the following iteration from Equation 22, 23 and 24:
$\mathbf{U}(t)=(I-\mathbf{\Delta}_{t})\cdot\mathbf{U}(t-1),$ (26)
$\mathbf{\Delta}_{t}=\begin{pmatrix}1-\frac{1+\beta k}{1+\gamma\eta_{t}}&\beta k&-\beta b\\ -\frac{\alpha k}{1+\gamma\eta_{t}}&\alpha k&-\alpha b\\ 0&0&0\end{pmatrix}.$ (27)
Based on the iteration, $\mathbf{U}(t)$ at the end of the $T$ epochs is then:
$\mathbf{U}(T)=\prod_{t=1}^{T}(I-\mathbf{\Delta}_{t})\cdot\mathbf{U}(0),$ (28)
where $\mathbf{U}(0)$ denotes $\begin{bmatrix}U_{v,0},&U_{s,0},&1\end{bmatrix}^{\top}$. The meaning of $U_{v,0},U_{s,0}$ is defined in Equation 24.
Under the circumstances, the following approximation holds true:
$\prod_{t=1}^{T}(I-\mathbf{\Delta}_{t})\approx\prod_{t=1}^{T}e^{-\mathbf{\Delta}_{t}}\approx e^{-\sum_{t=1}^{T}\mathbf{\Delta}_{t}}\approx e^{-\mathbf{\Delta^{\prime}}},$ (29)
$\mathbf{\Delta^{\prime}}=\begin{pmatrix}T-(1+\beta k)(T-\gamma\sum_{t=1}^{T}\eta_{t})&T\beta k&-T\beta b\\ -\alpha k(T-\gamma\sum_{t=1}^{T}\eta_{t})&T\alpha k&-T\alpha b\\ 0&0&0\end{pmatrix}.$ (30)
The first approximation is based on matrix Taylor expansion $e^{\mathbf{A}}=\sum_{k=0}^{\infty}\frac{\mathbf{A}^{k}}{k!},\lVert\mathbf{A}\rVert<1$, given matrix $\mathbf{\Delta}_{t}(t=1,\dots,T)$ is relatively small.
The second approximation is based on the fact that $\eta_{t}(t=1,\dots,T)$ is a relatively small quantity, considering that the total epoch $T$ is a large number. Therefore, the difference matrix between $\mathbf{\Delta}_{i}$ and $\mathbf{\Delta}_{j}(i\neq j,i,j=1,\dots,T)$ will be close to zero, making any two matrices $\mathbf{\Delta}_{i},\mathbf{\Delta}_{j}(i\neq j,i,j=1,\dots,T)$ approximately commutative.
The third approximation is based on the approximation $\frac{1}{1+\gamma\eta_{t}}\approx 1-\gamma\eta_{t}$, given $\eta_{t}$ a small quantity.
Equation 29 and 30 indicates that one can derive a solution of $U_{s}(T)$, which is only related to $\sum_{t=1}^{T}\eta_{t}$. Therefore, one can come to the conclusion that (i)solver uncertainty at the final epoch $T$ is approximately the same for cross-improvement with $\sum_{t=1}^{T}\eta_{t}=1$, and (ii)with the two elements in matrix $-\mathbf{\Delta^{\prime}}$ that $\gamma$ appears in both negatively correlated to the term $\gamma\sum_{t=1}^{T}\eta_{t}$, the final solver uncertainty is also negatively correlated to this term. This indicates that cross-improvement with $\gamma>0,\sum_{t=1}^{T}\eta_{t}=1$ outperforms self-improvement with $\gamma>0,\sum_{t=1}^{T}\eta_{t}=0$, in terms of solver capability.
Appendix C Review and Prospect
In this section, we discuss the difference between our work and prior works. We also discuss the challenge and perspective of self-improvement.
Different definition: In previous work [Song et al., 2025], the solver capability is defined as the accuracy, while the verifier capability is defined as the accuracy of responses that the model deems good responses. The calculation of accuracy depends on the golden answer contained in the original dataset, which is the most significant difference from our definition. In practice, many datasets do not contain gold answers, so our definition could be more widely applied.
Different concerns: Most of the prior works focus on finding a method that could make self-improvement more efficient. Instead, we pay attention to model the dynamics of self-improvement, which helps to understand the progress of self-improvement. In addition, we propose a new theoretical framework for cross-improvement which could be regarded as a solution to break through the limit of self-improvement.
Challenge: (i) Under supervised fine-tuning, self-improvement may be misled by incorrect responses. Thus, ensuring that the verifier outputs correct responses is a challenge. (ii) The current self-improvement method can only improve the model’s capability in similar tasks by optimizing a certain task. Finding a method to improve the performance of the model on different tasks is a challenge.
Future: (i) We propose a new framework for self-improvement, which provides a new perspective on self-improvement. Developing self-improvement algorithms based on our framework is a potential direction in the future. (ii) As there are no significant differences in the results of different allocation strategies, external data can be used to fine-tune a well-self-improved model to further improve the performance of the model. Compared with cross-improvement of the initial model, this method could reduce the resource consumption of training and obtain similar results.